OpenStudy (king_of_kings):

Evaluate. 7 × [8 + (16 – 9)] ÷ 5

OpenStudy (king_of_kings):

I NEED HELP

OpenStudy (logo):

70.2?

OpenStudy (king_of_kings):

are u shure

satellite73 (satellite73):

first \(16-9=?\)

OpenStudy (logo):

Oh right, do the boxes first

OpenStudy (king_of_kings):

7

OpenStudy (logo):

Brackets* so 7*15÷5

satellite73 (satellite73):

ok next it \(8+7=?\)

OpenStudy (king_of_kings):

15

satellite73 (satellite73):

then \(7\times 15=?\)

OpenStudy (king_of_kings):

85

satellite73 (satellite73):

oh hold on

satellite73 (satellite73):

\[7\times 15=?\]

OpenStudy (logo):

21 is right. That's what I got

OpenStudy (king_of_kings):

17

satellite73 (satellite73):

seven times fifteen \[\huge 7\times 15=?\]

OpenStudy (king_of_kings):

105

satellite73 (satellite73):

right, and finally \[105\div 5=?\]

OpenStudy (king_of_kings):

21

satellite73 (satellite73):

finished

OpenStudy (king_of_kings):

thnks so much

satellite73 (satellite73):

yw